OpenStudy (anonymous):

Ethylene glycol (C2H6O2) is used as an antifreeze in cars. If 400 g of ethylene glycol is added to 4.00 kg of water, what is the molality? Calculate how much the freezing point of water will be lowered. The freezing-point depression constant for water is Kf = –1.86°C/m. Show your work.

OpenStudy (anonymous):

molar mass of ethylene glycol = 2*(12.01078)+ 6(1.007947) + 2(15.99943)= 62.0 g mass of solvent water given = 4.0kg mass of ethylene glycol Given = 400g molality = 400/(62∗4) =1.61m with i =1, Kf = -1.86 C/m, and m = 1.61 delta Tf =i* Kf * m delta Tf =1* -1.86 * 1.61 =- 2.99C